jen boarded a bus bound for baguio at exactly 12AM. After one and a half hours, her boyfriend nomz followed driving the recently bought scooter at 80mph(miles per hour). Assuming the bus was travelling at an average speed of 60mph, at what time would Nomz overtake the bus?: jen boarded a bus bound for baguio at exactly 12AM. After one and a half hours, her boyfriend nomz followed driving the recently bought scooter at 80mph(miles per hour). Assuming the bus was travelling at an average speed of 60mph, at what time would Nomz overtake the bus?
Answer by stanbon(17981)

jen boarded a bus bound for baguio at exactly 12AM. After one and a half hours, her boyfriend nomz followed driving the recently bought scooter at 80mph(miles per hour). Assuming the bus was traveling at an average speed of 60mph, at what time would Nomz overtake the bus?
-----------------
Scooter DATA;
rate = 80 mph ; time = x hrs ; distance = 80x miles
-----------------
Bus DATA:
rate = 60 mph ; time = (x + 3/2) hr ; distance = 60(x+3/2) = 60x + 90 miles
----------------
EQUATION:
bus distance = scooter distance
60x + 90 = 80x
20x = 90
x = 9/2 = 4 1/2 hrs
-----------------------
They are together at 4:30 AM
=================
Cheers,
Stan H.